OpenStudy (anonymous):

50 = –6x + 4x – 8 solve?

OpenStudy (anonymous):

i got -15 and i right?

OpenStudy (anonymous):

x = -29

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

50 = -6x + 4x - 8 50 = -2x - 8 +8 +8 58 = -2x --------- -2 -2 -29 = x

OpenStudy (nurali):

50 = –6x + 4x – 8 Add both side by 8 50=-2x-8 50+8=-2x 58=-2x x=-58/2 x=-29

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

thank you for the explanation! :)

OpenStudy (nurali):

Anytime.